OpenStudy (anonymous):

[6.02] For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 9 −x − y = −8 −8 + y −8 − y 8 + y 8 − y

OpenStudy (anonymous):

To isolate the x in first equation, add 2y to both sides of the equal sign of the first equation x - 2y = 6 x - 2y + 2y = 6 + 2y x = 6 + 2y

OpenStudy (anonymous):

i dont see that in the choses the four choses are posted hun